Hello, this is the only information my teacher gave, please work on the problems that are possible with the available information, the one's that cannot be worked please note below the question. The teacher did not provide the balance sheet. Thank you.

Electricity Ltd is an expanding private company in the electric trade.

            Accounts prepared in January 1991 include the following information

Profit statement for the year ended 31st December

1989

1990

f

f

Sales

2000

3200

Less: cost of goods sold

1100

1800

Gross profit

900

1400

Less: Trading expense

       : Trading profit

450

450

550

850

Less: Debenture interest

25

25

Net profit before taxation

425

825

Less: Corporation tax

160

320

Net profit after taxation

265

505

Less: Ordinary share dividend

125

175

Undistributed profit for the year

140

330

    

Required;

Computation of the following accounting ratios;

  1. Return on capital
  2. Assets turnover
  3. Gross profit percentage
  4. Net profit percentage
  5. Current ratio
  6. Quick/acid test ratio.                                                                      (10 Marks)
Answer #1

Here it is possible to calculate only .

Gross profit ratio and net profit ratio

For all other ratios to be calculated we need balance sheet from where we can calculated assets, capital, debt, equity, current assets and current liability

so in the absence of balance sheet following ratios are not calculated

Return on capital

Assets turnover

Current ratio

Quick/acid test ratio.

Answer is as follow

Calculation of Gross profit ratio

Gross profit ratio = Gross Profit*100/Net Sales

Here Gross Profit =

For 1989 = f 900

1990 = f 1400

Sales are

For 1989 = f 2000

1990 = f 3200

So Gross profit ratio is as following

For 1989 = 900*100/2000 = 45%

For 1990 = 1400*100/3200 = 43.75%

Calculation of Net Profit Ratio

Net profit ratio = Net Profit*100/Net Sales

Here Net Profit after tax =

For 1989 = f 265

1990 = f 505

Sales are

For 1989 = f 2000

1990 = f 3200

so net profit ratio is as following

For 1989 =265*100/2000 = 13.25%

For 1990 = 505*100/3200 = 15.78% (rounded off to two decimals)   

FEEL FREE TO ASK ANY CLARIFICATION IF ANY REQUIRED KINDLY PROVIDE FEED BACK BY THUMBS UP IF SATISFIED IT WILL BE HIGHLY APPRECIATED

THANK YOU